OpenStudy (anonymous):

If the length of a rectangle is x + 4 and the width is 2x – 3, what is the area? A. 3x + 1 B. 6x + 2 C. 2x² + 5x – 12 D. 2x² – 5x – 12

OpenStudy (zehanz):

You have to multiply the two...

OpenStudy (zehanz):

So (x+4)(2x-3)=... Whatever it is, you can safely rule out options A and B. Why?

OpenStudy (anonymous):

ok i dont know but i might one day

OpenStudy (zehanz):

Do you know how to do (x+4)(2x-3)?

OpenStudy (zehanz):

Today is that day, btw...

OpenStudy (anonymous):

C

OpenStudy (zehanz):

@CR-Jennifer : giving a direct answer is not very helpful...

OpenStudy (zehanz):

YOu have to FOIL (x+4)(2x-3). FOIL means: First: x*2x = ... Outer: x*-3 = ... Inner: 4*2x = ... Last: 4*-3 = ...

OpenStudy (anonymous):

(x+4)(2x-3)=(x time 2x)+(x time (-3)+(4 time 2x) +(4 time(-3))=2x^2-3x+8x-12=2x^2+5x-12

OpenStudy (zehanz):

@brittanyburton: Do you understand now?